Lisa MacIntosh is the product manager for the Genius Smartwatch Manufacturing Company.   The company has total fixed costs of $6,090,000 for the coming year. The owner estimates that variable costs for this smartwatch will be $130. The company sells its product directly to selective retailers and the retail selling price will be $450. The Genius Company has priced its smartwatch so that the retailer will earn a 40% markup on the watch. The total market for this style of watch is currently 800,000 and Genius has a 15% market share.

A.         What price is the Genius Company charging its retailers for the watch? What is the manufacturer’s contribution per unit and the contribution margin percentage for each smartwatch sold to the retailer?   (3 points)

B.         How many units must be sold in order for Genius to break even? Is                              the smartwatch profitable? What is its profit or loss? (2 points)

C.         The Genius company's marketing director wants to generate profits

equal to 20% of sales for the watch line. How many units must be produced and sold in order to achieve this target return? What sales revenue must be generated in order to meet this target return? (2 points)

D.         Ms. MacIntosh is recommending that the advertising budget be

increased by $200,000. As a result of this increased spending on advertising, she expects to increase sales by 8%. Assuming that her forecast is correct, should the advertising budget be increased? What is the break even for the additional expenditure in units? What is the profit impact of such a move? (3 points)

E.         Ms. MacIntosh’s assistant just got a degree in economics and tells her that instead of increasing the advertising budget, she should decrease the price to the retailer by $20. He estimates that if she decreases the price to the retailer by $20, the company will be able to sell 8% more smartwatches to its retailers than she is currently selling. Mr. Wolf is considering the two alternative actions to increase profits: increase the advertising budget or decrease the price. Assuming her assistant’s forecast is correct, which action (increase advertising or decrease price) should Ms. MacIntosh recommend? What would the breakeven in units be after the decrease in price? What is the difference in profit between the two actions? (5 points)
Answer #1

Fixed cost = 6090000

Variable cost = 130

Selling price to customers = 450

Selling price to retailers = 450/(1+40%) = 321.4

Total market = 800000

Expected sales = 15% of 800000 = 120000

A.

Genius Company is charging retailers $321.4

Manufacturer contribution per unit is 321.4 – 130 = 191.4

The contribution margin percentage is 191.4/130 = 1.4723 or 147.23%

B.

BEP = Fixed cost/Contribution per unit

BEP = 6090000/191.4 = 31818.18

31819 units must be sold to break even. Considering that the market share is 15% and they can sell 120000 units, the business is profitable.

The profit is 120000*191.4 – 6090000 = 16878000

C.

To generate profit of 20% let the volume be X. This means that

191.4X – 6090000 = 0.2*(321.4X)

X = 47907.48

The company will have to sell 47908 units to make a profit of 20%

The sales revenue will be 321.4*47908 = 15397631.2

D.

Increase in advertising means increase in fixed cost

New fixed cost = 6090000 + 200000 = 6290000

New revenue = 321.4*120000*1.08 = 41601600

The new breakeven point is 6290000/191.4 = 32863.1 or 32864 units

Additional units = 32864-31819 = 1045 units.

The profit impact will be an additional 1045*191.4 = 200,013

Since 200013 is slightly more than 200000 cost, the advertising budget can be increased. The new breakeven point for the additional increase is 1045 units. The profit contribution will be $13